OpenStudy (anonymous):

A right cylinder has a diameter of 40 mm and a height of 30 mm, as shown, What is the volume of the cylinder? plz help

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Rembmer that V = lwh

OpenStudy (anonymous):

but i have two of those and not th ree

OpenStudy (compassionate):

Do you have an image attached to the problem?

OpenStudy (anonymous):

1 min

OpenStudy (compassionate):

To calculate the volume of a cylinder we need to know the radius of the circular cross-section of the cylinder - this is the measurement from the centre of the circle, to the outer-edge.

OpenStudy (compassionate):

To get the radius we take the diamater and divide it by 2. So 40/2 = 20 <-- That's our radius Now we need to multiply that by the height 33mm

OpenStudy (anonymous):

the height is 30 not 33

OpenStudy (compassionate):

Okay, we're measuring in milimeters, correct? So if I take 20mm and 30mm, then do the math, the answer comes out: 37714.28

OpenStudy (anonymous):

the answers are A. 37,680 mm3 B. 1256 mm3 C. 12,000 mm3 D. 150,720 mm3

OpenStudy (compassionate):

One minute. Let me recheck my math

OpenStudy (compassionate):

[Please wait... Verifyting answer...]

OpenStudy (compassionate):

\[V = \pi r^2h\] r = 20: \[20^2\] * 3.14 * 30

OpenStudy (compassionate):

So find twenty=squard, then multiply it by pi and 30.

OpenStudy (anonymous):

20 sqrt is 4.472135954999579

OpenStudy (compassionate):

Volume = (pi)r^2h = (pi)(.5d)^2h pi(20)^2(30) 12000(pi) 37699 mm^3

OpenStudy (anonymous):

so c

OpenStudy (compassionate):

A. ;)

OpenStudy (anonymous):

but a is A. 37,680 mm3

OpenStudy (compassionate):

I know ;) Choose it. It'll be correct.

OpenStudy (compassionate):

200%

OpenStudy (anonymous):

thank you bye